40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 243 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  4875   Fri Jun 24 01:05:32 2011 NicoleSummarySUSTransfer Function Model Analysis Summary and New Posted LED V vs. I Curve

I have updated the TT suspension wiki to include a new page on my transfer function model. In this new page, an introduction and analysis of my transfer function (including a comparison of the transfer functions for a flexibly- and rigidly-supported damper) are included.  This page contains linear and logarithmic bode plots.  Here is a link to the transfer function page.

 

I have also updated my photosensor page on the TT suspension wiki so that the experimental data points in my current versus voltage plot are plotted against the curve provided by the Hamamtsu data sheet. I have also included an introduction and analysis for my mini-experiment with the forward voltage and forward current of the LED. Here is link to the photsosensor page.

  4874   Fri Jun 24 00:13:24 2011 kiwamuUpdateABSLbeam profile measurement of LWE

The beam profile of the LWE (LightWave Electronics) NPRO was measured.

Mode matching telescopes will be designed and setup soon based on the result of the measurements.

 

Here is a plot of the measured beam profile.

beam_profile.png

 (some notes)

The measurement was done by using Kevin's power attenuation technique (#3030).

An window was put just after the NPRO and the reflected beam was sampled for the measurement to avoid the beam scan saturated.

  4873   Thu Jun 23 23:54:29 2011 KojiOmnistructureEnvironmentWe are saved

Sonali, Ishwita, and another anonymous SURF saved the long-lasted water shortage of the 40m

Attachment 1: IMG_0023.jpg
IMG_0023.jpg
  4872   Thu Jun 23 22:59:45 2011 kiwamuUpdateABSLI-P curve of LWE

 The I-P curve was measured again, but this time in a lower current range of 1.0-1.9 [A].

The plot below is the latest I-P curve.

IP_curve_small.png

(Decision)

Based on the measurement and some thoughts, I decided to run this laser at about 1.8 [A] which gives us a middle power of ~ 360 [mW].

In the 40m history, the laser had been driven at 2.4 [A] in years of approximately 2006-2009, so it's possible to run it at such a high power,

but on the other hand Steve suggested to run it with a smaller power such that the laser power doesn't degrade so fast.


(notes)

  The laser controller handed from PK (#4855) was used in this measurement.

The nominal current was tuned to be 1.8 [A] by tuning a potentiometer on the laser head (see page.18 on the manual of LWE).

There was a huge bump around 1.4 [A] and sudden power drop at 1.48 [A] although I don't know the reason.

Quote from #4842

The old days the NPRO ( inside the MOPA ) was running ~1.7A  500 mW

  4871   Thu Jun 23 22:53:02 2011 kiwamuUpdateCDSran activateDQ.py

I found some DQ channels (e.g. SENSOE_UL and etc.) for C1SUS haven't been activated, so I ran activateDQ.py.

Then I restarted daqd on fb as usual. So far the DQ channels look working fine.

  4870   Thu Jun 23 22:39:34 2011 JenneUpdatePSLPMC found unlocked

I found the PMC unlocked.  Koji noticed that the FSS Slow Actuator Adjust was railed at the positive end of the slider.  I set it close to zero, and relocked the PMC.  The FSS slow loop servo is doing its thing, and the PMC and MC are now locked. 

  4869   Thu Jun 23 22:00:22 2011 JamieUpdateSUSburt snapshot

I recorded a burt snapshot of these settings: /opt/rtcds/caltech/c1/burt/autoburt/snapshots/2011/Jun/23/21:40

  4868   Thu Jun 23 21:35:46 2011 Jamie, Rana, KiwamuUpdateSUSFix calibration for sus sensors

We have fixed the counts-to-micron (cts2um) calibration for the suspension sensor filters. Each suspension sensor filter bank (e.g. ULSEN) has a "cts2um" calibration filter. These have now been set with the following flat gains:

   40 V       10^3 um         um
 -------- *  --------  = .36  --
 2^16 cts     1.7 V           ct

The INMTRX was also fixed with proper element values:

UL UR LR LL SIDE  
.25 .25 .25 .25 0 POS
1.666 1.666 -1.666

-1.666

0 PIT
1.666 -1.666 -1.666 1.666 0 YAW
0 0 0 0 1 SIDE

This was done for all core optic suspensions (BS, PRM, SRM, ITMX, ITMY, ETMX, ETMY).

 

  4867   Thu Jun 23 21:34:21 2011 kiwamuUpdateCDSno foton on the CentOS machines

For some reasons foton's deafault sample rate is NOT correct when it runs on the CentOS machines.

It tries to setup the sample rate to be 2048 Hz instead of 16384 Hz until you specify the frequency.

To avoid an accidental change of the sample rate,

running foton on CentOS is forbidden until any further notifications.

Run foton only on Pianosa.

 

Additionally I added an alias sentence in cshrc.40m such that people can not run foron on CentOS (csh and tcsh, technically speaking).

Below is an example of raw output when I typed foron on a CentOs machine.

    rossa:caltech>foton
    DO NOT use foton on CentOS

  4866   Thu Jun 23 10:35:12 2011 steveUpdateComputersold computers leaving the lab

Rod Luna picked up these computers for Larry Wallace yesterday: Dell Inspiron 530, Dell Dimension 4600 and SunBlade 1000

Attachment 1: P1070905.JPG
P1070905.JPG
Attachment 2: P1070904.JPG
P1070904.JPG
Attachment 3: blade1000.JPG
blade1000.JPG
  4865   Thu Jun 23 10:17:49 2011 steveUpdatePSLAP table is open to PSL again

Access to the north side of the PSL table is blocked by the    8" beam guard. This opens the beam pathways between them.

Attachment 1: P1070911.JPG
P1070911.JPG
Attachment 2: P1070910.JPG
P1070910.JPG
  4864   Thu Jun 23 09:46:16 2011 ranaUpdateLSCPRMI locking : not stable enough

All the suspensions are bad until you fix them. But, ... there is a script which can be used to diagnose them today:

Python SUStest

  4863   Thu Jun 23 05:22:51 2011 kiwamuUpdateLSCPRMI locking : not stable enough

I was trying to measure the sensing matrix in the PRMI configuration, but basically gave up.

It is mainly because the lock of PRMI wasn't so stable and it didn't stay locked for more than a minute.

It looked like an angular motion fluctuated a lot around 1- 3 Hz. The beam spot on the AS camera moved a lot during the lock.

I have to figure out who is the bad suspension and why.

  4862   Thu Jun 23 02:12:12 2011 SonaliUpdateGreen Lockingwork schedule.

 June 22-June 24:

1.Coupling light into fibre at the ETMY.

2.Routing of the fibre to the PSL table.

June 27-June 30:

1.PSL optical table layout sketching.

2.Combining the PSL beam with fibre output onto a BS and then superpose them on a New Focus 1611 PD.

July 5-July 8:

1.Conversion of the PD output to voltage using MFD(Mixer Frequency Discriminator).

2. Report writing.

July 7: 5:00 pm: 1st Report Due.

July 11-July 22:

1.Locking Y-arm to PSL.

2.Setting up the feedback loop using the MFD output as the error signal and acting on the AUX laser frequency.

July 25-Aug 5:

1.Y-Arm cavity characterisation.

Measurement of the transmission of IR and green light through the cavity.

2.Analysis.

To obtain FSR, Finesse,Loss of the Cavity, Visibility, Transverse Modes(g-factor, astigmatism), Reflectivity, Q-factor.

3.Report and abstract writing.

Aug 1: 5:00 pm: 2nd Report and absract due.

Aug 8-11:

Preparation for talk and seminar.

  4861   Wed Jun 22 21:36:41 2011 ranaSummaryGeneralJuly 2011 vent plan

I put a paper Peet's bag with half of the Mini-Moos into George.

  4860   Wed Jun 22 18:51:47 2011 SureshUpdateIOOWFS2 RF response

I have shifted the Jenne laser back to the small table where we do RF PD characterisation (RFPD table).  I found several 25pin D-type connector cables, connected them in tandem and am using that to power the WFS2 sensor head at the RFPD table. 

The set up is ready for looking at the RF response of the  WFS sensors.  Will continue tonight.

 

  4859   Wed Jun 22 18:50:45 2011 JamieSummaryGeneralJuly 2011 vent plan

Kiwamu and I have started to put together a vent plan on the 40m wiki:

http://blue.ligo-wa.caltech.edu:8000/40m/vent/201107

We will keep working on this (there's still a *lot* to fill in), but please help fill in the plan by adding questions, answers, procedures, preparations, etc.

 

  4858   Wed Jun 22 18:41:23 2011 NicoleSummarySUSBROKEN bread board circuit box and L9337 LED Current Versus Voltage Curve

NOTE: The potentiometers on the bread board circuit box (the one I have been using with the signal generator, DC power, LED displays, and pulse switches) is BROKEN!

The potential across terminals 1 and 2 (also 2&3) fluctuates wildly and there dial does not affect the potential for the second potentiometer (the one with terminals 4, 5, and 6).

This has been confirmed by Koji and Jaimie.  PS I didn't break it! >____<

 

NEVERTHELESS, using individual resistors and the 500 ohm trim resistor, I have managed to get the current versus forward voltage plot for the Hamamatsu L9337 Infared LED

LED_I_vs_V_exp_plot.png

  4857   Wed Jun 22 17:42:03 2011 kiwamuUpdateLSCsensing matrix measurement

The sensing matrix was measured in the DRMI configuration for the first time.

The measurement was done by an automatic script and the realtime LOCKIN module built in the c1lsc model.

The resultant matrix is still too primitive, so I will do some further analysis.

 

(Measurement of sensing matrix)

 The quantities we want to measure are the transfer functions (TFs) from displacement (or change in optical phase) of each DOF to sensors in unit of [counts/m].

So essentially the measurement I did is the same as the usual TF measurement. The difference is that this measurement only takes TFs at a certain frequency, in this case 283 Hz.

 The measurement goes in the following order :

  (1) Lock DRMI

  (2) Shake an optic of interest longitudinally with an amplitude of 1000 counts at 283.103 Hz, where no prominent noise structures are present in any spectra of the sensor signals.

  (3) Put a notch filter at the same frequency of 283.103 Hz in each DOF (MICH, PRC and SRC) to avoid unwanted suppression due to the control loops.

       (This technique is essentially the same as this one, but this time the control loops are shut off only at a specific frequency )

       The notch filter I put has a depth of 60 dB and Q of 20. The filter eats the phase of ~10 deg at 200 Hz, which still allow servos to run with a high UGF up to 200Hz.

  (4) Take the output signal from a signal port of interest (i.e. REFL11_I, etc.,) and then put it into the realtime LOCKIN module.

  (5) Measure the resultant I and Q signals coming out from the LOCKIN module.

  (6) Repeat the procedure from (2) through (5) for each optic and sensor.

 

(Results)

 Again, the resultant sensing matrix is still primitive, for example the optic-basis should be converted into the DOF basis.

The values listed in the matrix below is the absolute values obtained by operation of sqrt( I^2 + Q^2) plus the polarity according to the output from I and Q of LOCKIN.

Therefore they still contain the actuator response, which is not desired. i will calibrate them into [counts/m] later by using the calibration factor of the actuator responses.

All the raw data showed the relative phase between I and Q either ~ 127 deg or ~ -53 deg.

In my definition, the one has 127 deg is plus polarity and the one has -53 deg is minus polarity.

Technically speaking the polarity depends on the polarity of the actuator and also the direction of the actuator against the DOFs.

Without any excitation the absolute values fluctuated at about 10-4 - 10-5, so the excitation amplitude was big enough to observe the sensing matrix.

Though, I still need to estimate the statistical errors to make sure the SNR is reasonably big.

 

sensemat.png

  Fig.1 Measured sensing matrix from optic to sensors.

 

(Things to be done)

  - convert the optic-basis (i.e. BS, ITMs, PRM and SRM) to the DOF-basis (i.e. MICH, PRC and SRC) so that the matrix is understandable from point of view of the interferometer control.

  - estimate the optimum demodulation phase for each DOF at each sensor port.

  - add some statistical flavors (e.g. error estimations and so on.)

  - edit the script such that it will keep watching the ADC overflows and the coherence to make sure the measurement goes well.

  - add some more signal ports (e.g. REFL55, POY55 and etc.)

  - compare with an Optickle model

Quote from #4851

The result will be updated later.

 

  4856   Wed Jun 22 17:35:35 2011 IshwitaUpdateGeneralHot air station

This is the new hot air station for the 40m lab.........

 

Attachment 1: P6220212.JPG
P6220212.JPG
Attachment 2: P6220213.JPG
P6220213.JPG
  4855   Wed Jun 22 15:24:10 2011 kiwamuUpdateABSLgot a laser controller for LightWave

Peter King came over to the 40m with a laser controller and gave it to us.

We will test it out with the LightWave NPRO, which was used for MOPA.

Attachment 1: DSC_3150.JPG
DSC_3150.JPG
Attachment 2: DSC_3153.JPG
DSC_3153.JPG
  4854   Wed Jun 22 12:29:57 2011 IshwitaSummaryAdaptive FilteringWeekly summary

I started on the 16th with a very intense lab tour & was fed with a large amount of data (I can't guarantee that I remember everything....)

Then... did some (not much) reading on filters since I'm dealing with seismic noise cancellation this summer with Jenne at the 40m lab.

I'll be using the Streckeisen STS-2 seismometers & I need to use the anti aliasing filter board that has the 4 pin lemo connectors with the seismometers & its boxes that require BNC connectors. I spent most of the time trying to solder the wires properly into the connectors. I was very slow in this as this is the first time I'm soldering anything.... & till now I've soldered 59 wires in the BNC connectors....

 

 

  4853   Wed Jun 22 12:24:44 2011 NicoleSummarySUSMidweek 2 Work Summary

I have made my transfer function model and posted it to the suspension wiki. Here is the link to my model!

Bode Plot Model

Please let me know if there need to be any adjustments, but I have posted the bode plots, a model image, and an explanation of why I think it's right! ^ ___^ V

I am currently working on the photo sensor circuit for the displacement detector. So far, I have gotten the infared LED to light up! ^ ___^ V

I am now trying to get a plot of forward voltage versus current for the LED. HOPEFULLY it will match the curve provided in the LED datasheet.

I'm using the bread board circuit box and when I'm not working at the bench, I have signs posted. PLEASE DO NOT REMOVE THE CONNECTIONS! It is

fine to move the bread board circuit box, but please do not disturb the connections > ____<

Here is a photo of the workspace

P6220200.JPG

  4852   Wed Jun 22 01:59:43 2011 SonaliUpdateGreen Lockingfibre-coupling of the IR beam

 What I did today.

1. Collimation of a beam.

  • I then practiced collimation of a 700 nm laser (output) beam after being coupled through a fibre.
  • I put together the set-up as shown in the attached picture where I used ....... to couple 650nm light into the PM.... fiber.
  • I kept shifting the focus of the output beam to an appreciable distance till it was approximately collimated.

2. Coupling of the IR light at the ETMY table to a fibre.

  • The fibre coupler was put in place to couple light into the fiber.
  • I put in the mirrors as planned to direct the IR beam exiting the doubling crystal towards the fiber coupler (input).
  • The mirrors were aligned such that the beam falls on the input lens of the coupler.
  • The far-end of the fibre originally would have gone to to PSL table but it has been put on this table to study the power of the IR beam transmitted through this set-up.  The output end of the fiber has been connected to another fiber optic coupler to collimate the exiting beam.
  • The picture of the current status attached.
Attachment 1: ETMY_june_21.jpg
ETMY_june_21.jpg
Attachment 2: collimation_700nm_21_june.JPG
collimation_700nm_21_june.JPG
  4851   Tue Jun 21 23:29:41 2011 kiwamuUpdateLSCsensing matrix measurement

I am now measuring the sensing matrix in the DRMI configuration.

A goal of tonight is to measure the sensing matrix as a test of the script.

 

The result will be updated later.

  4850   Tue Jun 21 20:35:50 2011 kiwamuUpdateLSCa script to measure sensing matrix

Last night I was making a script which will measure the sensing matrix using the realtime LOCKIN module.

The script is a kind of expansion of Jamie's one, which measure the asymmetry, to more generic purpose.

It will shake a suspended optic of interest and measure the response of each sensor by observing the demodulated I and Q signals from the LOCKIN module.

I will continue working on this.

 

  (current status)

 - made a function that drives the LOCKIN oscillator and get the data from the I and Q outputs.

 - checked the function with the MICH configuration.

   ITMX, ITMY and BS were shaken at 100 Hz and at different time.

   Then the response of AS55_Q showed agreement with what I got before for the actuator calibration (see this entry).

   It means the function is working fine.

Attachment 1: elog.png
elog.png
  4849   Tue Jun 21 19:54:33 2011 SureshUpdateGreen LockingLightWave NPRO power supply shifted to ETMY end table

The Lightwave NPRO power supply which is being shared between the AS table and the ETMY table has been shifted back to the ETMY table

The current to the laser is set at 1.5A.  The laser output is 200mW at this current level.

  4848   Tue Jun 21 19:10:29 2011 SureshUpdateEnvironmentETMY end table

[Suresh, Sonali]

There were small pieces of glass, remnants of a fluorescent tube, which were lying around on the ETMY end table for a while now.  We picked up the larger pieces by hand and used the HEPA filtered vacuum cleaner to pick up the remaining glass and dust on the table. 

  4847   Tue Jun 21 16:11:13 2011 steveUpdateSAFETYsurf safety training

40m surfs:  Nicole Ing, Iswita Saikia and Sonali Mohapatra received 40m specific safety training  today.

Attachment 1: P1070902.JPG
P1070902.JPG
  4846   Tue Jun 21 00:38:21 2011 SonaliUpdateGreen Lockingrepositioned "QPDY_PD"

1.The aim is the laser frequency stabilisation of PSL and AUX.

2.As a first step we want to couple some of the AUX laser beam into a single mode optical fibre and route the fibre to the PSL table.

3.The position of the optical fibre on the ETMY table is shown by the coupler in the attached picture. The yellow lines show the new scheme we want to implement.

4.WHAT WE DID TODAY.

  • The Y-arm was locked so that we could use the transmitted IR beam as the reference.
  • We shifted the position of the "QPDY_PD" .
  • We also shifted the "ETMYT" camera to make space for the "QPDY_PD".
  • The mirror  directing the beam into the "QPDY_PD" was rotated by 90 degrees to adhere to the new position of the "QPDY_PD".
  •  The attached photo shows the table as it is right now after the repositioning.
5.We continue with the positioning of the fibre-coupling tomorrow.
Attachment 1: ETMY_table_2011_June20.png
ETMY_table_2011_June20.png
Attachment 2: ETMY_table_change1.jpg
ETMY_table_change1.jpg
  4845   Mon Jun 20 18:36:49 2011 SureshUpdateLSCREFL55 PD update

[Suresh, Koji]

   I used a matlab code written by Koji to analyse the transimpedance and current noise data  of REFL55.  The details are in the attached pdf file.

Resonance is at 55.28 MHz:

Q of 4.5, Transimpedance of 615 Ohms

shot noise intercept current = 1.59 mA

current noise =21 pA/rtHz

 

Notch at 110.78 MHz:

Q of 54.8 Transimpedance of 14.68 Ohms.

 

 

 

 

 

 

 

 

 

Quote:

[Rana, Koji]

REFL55 was modified. The noise level confirmed. The PD is now ready to be installed.

 


Kevin's measurement report told us that something was wrong with REFL55 PD. The transimpedance looked OK, but the noise level was terrible (equivalent to the shotnoise of 14mA DC current).

Rana and I looked at the circuit, and cleaned up the circuit, by removing unnecessary 11MHz notch, 1k shunt resister, and so on.

I made a quick characterization of the PD.

First page:

The transimpedance ws measured as a function of the frequency. The resonance was tuned at 55MHz. The notch was tuned at 110MHz in order to reject the second harmonics. The transimpedance was ~540V/A at 55MHz. (For the calibration, I believed the DC transimpedance of 50V/A and 10000V/A for the DC paths of this PD and #1611, respectively, as well as the RF impedance (700V/A0 of #1611.

Second page:

Output noise levels were measured with various amount of photocurrent using white light from a light bulb. The measurement was perforemed well above the noise level of the measurement instruments.

Third page:

The measured output noise levels were converted into the equivalent current noise on the PD. The dark noise level agrees with the shot noise level of 1.5mA (i.e. 22pA/rtHz). In deed, the noise level went up x~1.5 when the photocurrent is ~1.4mA.

 

Attachment 1: REFL55_response.pdf
REFL55_response.pdf
  4844   Mon Jun 20 18:12:20 2011 NicoleUpdateSUSSmall Table Cleaned and Levelled

P6220198.JPG

The small optical bench (next to the MC-2 Chamber and the tool box tower) has been cleared of the misc. object previously on it, cleaned, and leveled (after much calibration X___X).

PLEASE, PLEASE, PLEASE do NOT MOVE OR HIT THE TABLE! It was incredibly painful to level.

This is how leveling the table made me feel...

P6220199.JPG

VERY SAD...so do not move please!

The shaker has already been moved to the table and the amplifier for my shaking experiment is located behind the table (not on the table, as to prevent scratching).

 

 

  4843   Mon Jun 20 17:58:00 2011 ranaUpdateCDSGateway program killed

There was a rogue, undocumented, gateway process running on NODUS since ~4 PM. This guy was broadcasting channels back into the Martian and causing lockups in the IOO controls. I did a kill -9 on its process.

Someone will pay for this.

  4842   Mon Jun 20 16:44:02 2011 steveUpdateABSLI-P curve of LightWave M126-1064-700

 

 Put the serial numbers into the elog. So we can identify the laser and controller in the future.

The old days the NPRO ( inside the MOPA ) was running ~1.7A  500 mW

  4841   Mon Jun 20 13:48:25 2011 KojiUpdateABSLI-P curve of LightWave M126-1064-700

Hmm. Was the current within the operating range? (i.e. Is it a 700mW laser or a 1W one?)

You can obtain the nominal operating current from the old EPICS values (or some elog entries).

Note that NPROs are designed to be healthy only at around the nominal pumping power
(i.e. thermal gradient, and thermal lensing of the crystal, etc.)

ALSO:

Be aware that this laser should be used under the old SOP. So the appropriate interlocking is mandatory.

And probably we need to modify the SOP such that it reflects the latest situation.

Quote:

The I-P curve of the LightWave NPRO, which was taken out from the MOPA box, was measured. It looks healthy.

The output power can go up to about 1 W, but I guess we don't want it to run at a high power to avoid any further degradation since the laser is old.

  4840   Mon Jun 20 11:38:49 2011 kiwamuUpdateABSLI-P curve of LightWave M126-1064-700

The I-P curve of the LightWave NPRO (M126-1064-700), which was taken out from the MOPA box, was measured. It looks healthy.

The output power can go up to about 1 W, but I guess we don't want it to run at a high power to avoid any further degradation since the laser is old.

 

IP_curve.png

 X-axis is the current read from the display of the controller. Y-axis is the output power, directly measured by Coherent PM10.

The measurement was done by changing the current from the controller.

Quote from #4832

 [Things to be done]

   - measure the beam profiles and power

   - get a laser controller, which will be dedicated for this laser, from Peter King

  4839   Mon Jun 20 11:04:03 2011 NicoleUpdateSUSWork Plan for Week 2

Here is my work plan for this week:

Current Week Plan (Week 2) (As of 6/17/11)

 

Setting Up for Horizontal Displacement Measurements

1) Help Steve clean small table for experiment

2) Remove aluminum base from TT suspension

3) Mount shaker onto table base

4) Mount horizontal slider onto table base

5) Connect TT suspension, shaker, and horizontal slider

Begin Assembly of Sensors

1) Begin building circuit for displacement photosensors

2) Calibrate photosensor using linear regions of power versus distance curves

3) Circuit box for photosensors?

  4838   Mon Jun 20 10:45:43 2011 JamieUpdateCDSPower restored to 1X1/1X2 racks. IOO binary output module installed.

All power has been restored to the 1X1 and 1X2 racks.  The modecleaner is locked again.

I have also hooked up the binary output module in 1X2, which was never actually powered.  This controls the whitening filters for MC WFS.  Still needs to be tested.

  4837   Mon Jun 20 09:28:19 2011 JamieUpdateCDSShutting down low-voltage DC power in 1X1/1X2 racks

In order to install the BO module in 1X2, I need to shut down all DC power to the 1X1 and 1X2 racks.

  4836   Mon Jun 20 09:04:13 2011 steveUpdateSUSETMX damping restored

ETMX sus damping restored

  4835   Mon Jun 20 00:59:02 2011 kiwamuSummaryGeneralWeekly report
This is a summary for the week ending June 19th. Feel free to edit this entry.
(Number of elog entries = 27)

* Refinement of LSC screen
    -> Kissel buttons and some indicators were newly installed
    -> A script to autonatically generate kissel buttons was made

* New BIO installed on ETMY

* LightWave for ABSL
    -> taken out from the MOPA box and put on the AP table with temporary use of the Y end laser controller
 
* Shipping 2 RFPDs to LLO
 
* LEDs on the BIO for the vertex suspensions were blown
    -> fixed and re-installed. A test script will be prepared
 
* PEM AA board was fixed
 
* A plot of the MICH noise was produced for the first time
 
* Schnupp asymmetry measurement : Las = 3.64+/-0.32cm
 
* The photo diode on WFS2 has been replaced by YAG-444-4A
 
* SUS binary IO crates were taken out
 
* Fiber died
     ->C1LSC was unable to communicate to C1SUS. Installing a new copper Dolphine fixed the issue.
 
* SURF students came
  4834   Fri Jun 17 23:20:05 2011 KojiUpdateLSCSome updates of the LSC screen

Some updates of the LSC screen

- Signal amplitude monitor for the PD signals (--> glows red for more than 1000)

- Kissel Buttons for the main matrices

- Trigger display at the output of the DOF filters

- Signal amplitude monitor for the SUS LSC output (--> glows red for more than 10000)

 

ADC Over flow monitor is showing some unknown numbers (as ADCs are handled by IOPs).
I asked Joe for the investigation (and consideration for the policies)

Attachment 1: screen.png
screen.png
  4833   Fri Jun 17 17:02:15 2011 JamieUpdateSUSETMX/ETMY binary output modules (re)installed, not yet tested

I have installed a new binary output module in ETMY, where there was none previously.  It is installed, powered (with working LEDs), hooked up (to the binary output card and the cross connect), but it hasn't been fully tested yet.

I also re-installed the binary output module in ETMX, with newly modified power-indicator LEDs.

Both modules are fully installed, but they have not yet been fully tested to confirm that they are indeed switching the whitening and de-whitening filters.

  4832   Fri Jun 17 16:05:07 2011 kiwamuUpdateABSLLightWave out of MOPA box

[Suresh / Kiwamu]

 We did the following things :

   - Took the LightWave NPRO out from the MOPA box

   - Temporarily took out the laser controller which has been connected to the Y end laser.

   - Put the LightWave on AP table and plugged the laser controller and confirmed that it still emits a beam

 DSC_3139_small.png

 

[Things to be done]

   - measure the beam profiles and power

   - get a laser controller, which will be dedicated for this laser, from Peter King

 

[Background and Motivation]

 The PRC and SRC length have to be precisely measured before the vent.

In order to measure those absolute length we are going to use the Stochino technique, which requires another laser to scan the cavity profiles.

The LightWave NPRO laser in the MOPA box was chosen for the Stochino laser because it has a large PZT range of 5 MHz/V and hence allows us to measure a wider frequency range.

The laser in the MOPA box had been connected to home-made circuits, which are not handy to play with. So we decided to use the laser with the usual laser controller.

Peter King said he has a LightWave laser controller and he can hand it to us.

Until we get the controller from him we do some preparations with temporary use of the Y end laser controller.

  4831   Fri Jun 17 08:03:48 2011 steveUpdateSUSITMY sus damping restored

ITMY sus damping restored.

 

  4830   Fri Jun 17 00:17:26 2011 ranaConfigurationElectronics2 RFPDs sent to LLO

Koji and I found 2 RFPD boxes to send to LLO. We've put them onto Steve's desk to be overnighted to Valera.

One of them is our old 21.5 MHz gold box RFPD from the FSS (which we don't use). The other one is a 2mm gold box one which was previously tuned for 66 MHz.

 

They shipped out on Friday

Attachment 1: P1070895.JPG
P1070895.JPG
  4829   Thu Jun 16 23:19:09 2011 KojiUpdateSUSVertex SUS Binary Output Boxes removed

[Jamie, Koji]

- We found the reason why some of the LEDs had no light. It was because the LEDs were blown as they were directly connected to the power supply.
The LEDs are presumably designed to be connected to a 5V supply (with internal current-limiting resistor of ~500Ohm). The too much current
with the 15V (~30mA) made the LED blown, or the life-time of them shorter.

- Jamie removed all of the BO modules and I put 800Ohm additional resister such that the resultant current is to be 12mA.
The LEDs were tested and are fine now.

- The four BO boxes for C1SUS were restored on the rack. I personally got confused what should be connected where
even though I had labeled for BO0 and BO1. I just have connected CH1-16 for BO0. The power supplies have been connected only to BO0 and BO1.

- I tested the whitening of PRM UL sensor by exciting PRM UL sensor. The transfer function told us that the pendulum response can be seen
up to 10-15Hz. When the whitening is on, I could see the change of the transfer function in that freq band. This is good.
So the main reason why I could not see theis was that the power supply for the BOs were not turned on.

- I suppose Jamie/Joe will restore all of the BO boxes on the racks tomorrow. I am going to make a test script for checking the PD whitenings.

  4828   Thu Jun 16 08:45:14 2011 steveUpdateSUSVertex SUS Binary Output Boxes removed

Quote:

- I was investigating the SUS whitening issue.

- I could not find any suspension which can handle the input whitening switch correctly.

- I went to 1X5 rack and found that both of the two binary output boxes were turned off.
As far as I know they are pulling up the lines which are switched by the open collector outputs.

- I tried to turn on the switch. Immediately I noticed the power lamps did not work. So I need an isolated setup to investigate the situation.

- The cables are labelled. I will ask steve to remove the boxes from the rack.

 I shut down damping to the Vertex optics and removed Binary IO  Adapter chassy BO0 and BO1

About a week ago I discussed the BO0's power indicator lights with Kiwamu. They were  not on or they were blinking on-off.

I put screws into ps connectors in the back, but it did not helped.

Attachment 1: P1070894.JPG
P1070894.JPG
  4827   Thu Jun 16 00:43:36 2011 KojiUpdateSUSVertex SUS Binary Output Boxes were turned off / need investigation

- I was investigating the SUS whitening issue.

- I could not find any suspension which can handle the input whitening switch correctly.

- I went to 1X5 rack and found that both of the two binary output boxes were turned off.
As far as I know they are pulling up the lines which are switched by the open collector outputs.

- I tried to turn on the switch. Immediately I noticed the power lamps did not work. So I need an isolated setup to investigate the situation.

- The cables are labelled. I will ask steve to remove the boxes from the rack.

  4826   Thu Jun 16 00:39:08 2011 KojiUpdatePEMPEM AA Board has been diagnosed and fixed

As seen in the photo, the board has a strange bulge on the board,
and
the color of the internal line around the bulge got darkened.

I don't trust this board any more. We should switch to the alternative one.

Quote:

Steve will re-install the board in the rack in the morning.

 

ELOG V3.1.3-